OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

3x+3=-2x+8 3x=-2x+5 5x=5 x=1

OpenStudy (anonymous):

All he did was multiply what is outside the bracket, with what is inside.

OpenStudy (anonymous):

Then solved for x.

hero (hero):

Here's a more interesting way to solve it: 3(x+1) = -2(x-1) + 6 3(x+1) - 6 = -2(x-1) 3((x + 1) - 2) = -2(x-1) 3(x-1) = -2(x-1) 3(x-1) + 2(x-1) = 0 5(x-1) = 0 x - 1 = 0 x = 1

hero (hero):

I suspect @jiji501 has doubts about my solution

OpenStudy (anonymous):

Or wait... yeah... I don't get it x_x

OpenStudy (anonymous):

@Hero Care to explain? :)

hero (hero):

The steps are plain as day. What exactly is it that you would like me to explain. I suppose you want me to explain the part where I factored out 3. or perhaps the part where I removed the parentheses from x + 1 , then added -2 to get x - 1.

OpenStudy (anonymous):

Oh... explain to me how birds can fly.